OpenStudy (anonymous):

Graph the following conditions. {(x, y): x + 2y < 6}

OpenStudy (anonymous):

ok so the little"<" thing that a line below it

OpenStudy (anonymous):

solve for y and plot normally then...

OpenStudy (anonymous):

the answers i have are in pics

OpenStudy (anonymous):

y <= 3-(1/2)*x choose a point for x any point like 0 then y=< 3 therefore shade below line make the line solid

OpenStudy (anonymous):

https://www.google.com/search?rlz=1C1CHFX_enUS470US470&aq=f&sugexp=chrome,mod=18&sourceid=chrome&ie=UTF-8&q=3-(1%2F2)*x looks like that with the shading under the line

OpenStudy (anonymous):

This site is really helpful https://www.desmos.com/calculator

OpenStudy (callisto):

First draw the line for x+2y = 6. Since it's with less than or equal to, the line is a solid line. To draw the line, Put x =0, 0 + 2y = 6 => y=3 So, (0,3) is one of the points lying on the straight line. Put y = 0 x + 2(0) =6 => x=6 So, (6, 0) is another point lying on on the straight line. Now, draw x-axis and y-axis on a piece of graph paper. Mark the scale. Plot the 2 points on the graph. Draw a straight line joining the two points. Write the equation (x+2y=6) near the line. To determine the shaded region, put a point to test. For instance, consider the point (0, 0) LS = 0 + 2(0) = 0 < 6 So, it satisfies the condition. That means the point (0, 0) is in the shaded region. Then, shade the region where (0, 0) is located at.

OpenStudy (anonymous):

how would teh graph look like?

OpenStudy (callisto):

Have you drawn it?

OpenStudy (anonymous):

^^ not in my answers

OpenStudy (callisto):

If you draw it, you'll know the answer.

OpenStudy (anonymous):

oh than like the google link i gave with it shaded above :P

OpenStudy (callisto):

I've mentioned it quite clearly in my previous post. Which one you don't know?

OpenStudy (anonymous):

sorry i got it thanks:D

OpenStudy (callisto):

I just quoted from my previous post. To draw the line, Put x =0, 0 + 2y = 6 => y=3 So, (0,3) is one of the points lying on the straight line. Put y = 0 x + 2(0) =6 => x=6 So, (6, 0) is another point lying on on the straight line.

OpenStudy (anonymous):

sorry i got it thanks:D

OpenStudy (callisto):

Welcome